OpenStudy (anonymous):

what is the equation for a line that passes through the point (6,-1) and is perpendicular to the line y + 2x = 5

OpenStudy (anonymous):

show work plzz

OpenStudy (aravindg):

line perpendicular to y+2x=5 is 2y-x=p now it passes through (6,-1) so 2*-1-6=p means p=-8 so eqn of line is 2y-x=-8

OpenStudy (anonymous):